# Meta-automatic Sequences

### v1

@@ 4 Meta(2)
Define $\mathcal{M}_{2}:\mathbb{N}_{0}\to\{0,1\}$ by $\mathcal{M}_{2}(0)=0$, $\mathcal{M}_{2}(1)=1$, and for $n\geq 0$, so that
$\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n)),\ \mathcal{M}_{2}(4n+1)=1-\mathcal{M}_{2}(4n)$ (12)
and
$\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-\mathcal{M}_{2}(n)),\ \mathcal{M}_{2}(4n+3)=1-\mathcal{M}_{2}(4n+2).$ (13)
Lemma 27.
@@
Well-definedness follows by induction as in Lemma 18, since the selector indices satisfy $2n+\mathcal{M}_{2}(n)\leq 2n+1<4n$ and $2n+1-\mathcal{M}_{2}(n)\leq 2n+1<4n+2$. Balancedness is immediate from the complementary definitions in (12)–(13). ∎
4.1 Lifted state and affine closure
Define the lifted state
@@ 4.1 Lifted state and affine closure
We use Definition 26, Lemma 28, and balancedness throughout. Write $s(n)=(x,y)$ so that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$.
Case $r=0$. The first coordinate is given by $\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n+x)$. By the recovery lemma, we have that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$. In both cases, for $x=0$ or $x=1$, a direct check gives that $\mathcal{M}_{2}(2n+x)=y$. For the second coordinate $\mathcal{M}_{2}(8n+y)$, we see that if $y=0$, then $\mathcal{M}_{2}(8n)=\mathcal{M}_{2}(4(2n))=\mathcal{M}_{2}(2\cdot 2n+\mathcal{M}_{2}(2n))$, and, since $\mathcal{M}_{2}(2n)=x\oplus y=x$, this equals $\mathcal{M}_{2}(4n+x)$. Simplifying both cases gives $x$. If $y=1$, then $\mathcal{M}_{2}(8n+1)=1-\mathcal{M}_{2}(8n)$, and, by the same analysis as before, we get $x\oplus 1$. In both sub-cases the second coordinate equals $x\oplus y$. Therefore, we obtain
@@
For the second coordinate of $s(4n+1)$, by definition, this is $\mathcal{M}_{2}(2(4n+1)+\mathcal{M}_{2}(4n+1))=\mathcal{M}_{2}(8n+2+(y\oplus 1))$. When $y=0$, we need $\mathcal{M}_{2}(8n+3)=\mathcal{M}_{2}(4(2n)+3)=1-\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$. Since $\mathcal{M}_{2}(2n)=x$, this equals $1-\mathcal{M}_{2}(4n+1-x)=x\oplus y$ for both values of $x$. When $y=1$, we need $\mathcal{M}_{2}(8n+2)=\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$. Since $\mathcal{M}_{2}(2n)=x\oplus 1$, this equals $\mathcal{M}_{2}(4n+x)=x\oplus y$ for both values of $x$. The second coordinate is therefore $x\oplus y$ in both cases. Therefore
@@
Case $r=2$. For the first coordinate $\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-x)$, by XOR linearization (Lemma 10), this equals $\mathcal{M}_{2}(2n+1)\oplus x=(x\oplus y\oplus 1)\oplus x=y\oplus 1$.
@@
The four letters of the lifted alphabet correspond to the states $(x,y)\in\mathbb{F}_{2}^{2}$ of the affine automaton (Theorem 29) via
$\mathsf{0}=(0,0),\quad\mathsf{1}=(0,1),\quad\mathsf{2}=(1,0),\quad\mathsf{3}=(1,1).$
With this identification, the morphism $\sigma_{\mathcal{M}_{2}}$ is derived directly from the affine rule $s\mapsto As+\mathbf{v}(d)$, since the image $\sigma_{\mathcal{M}_{2}}(\mathsf{r})$ lists the four successive states visited when the current state is $\mathsf{r}$, for digits $d\in\{0,1,2,3\}$ processed in order. The coding $\tau(\mathsf{r})=x_{\mathsf{r}}\bmod 2$ (first coordinate of the state) then recovers $\mathcal{M}_{2}$. Under the symmetric labeling $a=\mathsf{0}$, $b=\mathsf{1}$, $c=\mathsf{2}$, $d=\mathsf{3}$, the morphism reads $a\to acdb$, $b\to dbac$, $c\to bdca$, $d\to cabd$, with $\tau(a)=\tau(b)=0$ and $\tau(c)=\tau(d)=1$. This alternative notation—where the alphabet labels match the parity of the binary output—is sometimes more convenient for right-special factor analysis.
@@ 4.2 An explicit digit formula
With the notation above, we have that $\mathcal{M}_{2}(n)$ satisfies
@@
The mapping $s\mapsto As+\mathbf{v}(d)$ with $A^{3}=I$ (see Theorem 29 and Remark 30) gives
$s(n)=\sum_{j=0}^{L-1}A^{L-1-j}\,\mathbf{v}(d_{j})\pmod{2}$
after starting from $s(0)=(0,0)$ and processing $L$ base-44 digits MSB-first. Since $A^{3}=I$, the exponent reduces modulo 33. Computing $A^{k}\mathbf{v}(d)$ for $k\in\{0,1,2\}$ and each digit $d$, we obtain
@@
$q(n):=\sum_{\begin{subarray}{c}m\geq 0\\ m\ (\mathrm{mod}\ 6)\in\{0,1,3,4\}\end{subarray}}b_{m}2^{m},$ (15)
so that $q(n)$ is obtained from $n$ by zeroing the bits in positions $m\equiv 2,5\pmod{6}$. Then
@@
In Theorem 34, each base-$4$ digit $d_{k}$ contributes (via $i_{k},j_{k}$) to $\mathcal{M}_{2}(n)$ in a way that depends only on $k\bmod 3$: For $k\equiv 0$, we take $i_{k}\oplus j_{k}$, for $k\equiv 1$ we take only $j_{k}$, and for $k\equiv 2$ we take only $i_{k}$. Since $i_{k}$ and $j_{k}$ are precisely the binary digits $b_{2k}$ and $b_{2k+1}$, this means that $\mathcal{M}_{2}(n)$ is the XOR of all binary digits of $n$ except those in positions $2k$ with $k\equiv 1\pmod{3}$ and $2k+1$ with $k\equiv 2\pmod{3}$, i.e., except $m\equiv 2,5\pmod{6}$. Thus $\mathcal{M}_{2}(n)$ is the parity of the binary digit-sum of $q(n)$ defined in (15), hence $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$. The base-$4$ description of $q(n)$ is the same statement grouped into pairs of bits. ∎
@@ 5 Factor complexity
Let $p_{a}(n)$ denote the number of distinct length-$n$ factors in the infinite word $a(0)a(1)a(2)\cdots$.
@@
For every infinite binary word $a$ and every $n\geq 1$,
$p_{a}(n+1)-p_{a}(n)=\#\mathrm{RS}_{a}(n),$
where $\mathrm{RS}_{a}(n)$ denotes the set of length-$n$ factors $u$ of $a$ such that both $u0$ and $u1$ are factors of $a$.
@@
Lemma 37 (Regularity of the right-special count).
If $a$ is $k$-automatic, then the function $n\mapsto\#\mathrm{RS}_{a}(n)$ is $k$-regular in the sense of Allouche–Shallit [3], hence effectively computable from any DFAO for $a$. When $\mathrm{RS}_{a}(n)$ is additionally bounded, it is $k$-automatic.
@@
The predicate “$u$ is a length-$n$ factor of aa” and the predicate “$u$ is right-special” are first-order definable over $(\mathbb{N},+,V_{k})$ relative to the DFAO for $a$, so the set of encodings of such pairs is regular and recognized by a finite automaton. Counting right-special factors of a given length reduces to a finite-automaton computation, giving a $k$-regular sequence. For details and background, we again refer to the appropriate texts from Allouche and Shallit [3] and from Shallit [27]. If the values are bounded, $k$-regularity reduces to $k$-automaticity. ∎
Several claims in this section can be mechanically certified using the Walnut theorem prover [27], which decides first-order properties of automatic sequences in Presburger arithmetic. For clarity, and in the hope of developing a deeper understanding of the behavior of meta-automatic sequences, we present “non-Walnut” proofs wherever appropriate.
5.1 Quarto
We first record a key property of the Thue–Morse factor language.
@@ 5.1 Quarto
The language $\mathrm{Fac}(\mathbf{t})$ is closed under bitwise complementation, i.e., if $w\in\mathrm{Fac}(\mathbf{t})$, then $\bar{w}\in\mathrm{Fac}(\mathbf{t})$.
@@
By Proposition 16, on each dyadic block $2^{k},2^{k+1})$ the sequence $\mathcal{Q}$ coincides with $\mathbf{t}$ or $1-\mathbf{t}$ according to the parity of $k$. A factor of $\mathcal{Q}$ of length $n$ either lies entirely within one block or straddles a dyadic boundary.
Intra-block factors. If the factor lies in an even block it is a factor of $\mathbf{t}$. If it lies in an odd block it is a complement of a factor of $\mathbf{t}$, which by Lemma 38 is again a factor of $\mathbf{t}$. So all intra-block factors belong to $\mathrm{Fac}(\mathbf{t})$.
Junction factors. A factor straddling boundary $2^{k}$ has the form $u\cdot\bar{v}$ (or $\bar{u}\cdot v$, by a symmetric argument), where $u\cdot v\in\mathrm{Fac}(\mathbf{t})$. Such a word need not lie in $\mathrm{Fac}(\mathbf{t})$: If $u\cdot\bar{v}\in\mathrm{Fac}(\mathbf{t})$, then $\bar{u}\cdot v\in\mathrm{Fac}(\mathbf{t})$ by Lemma 38, but this need not hold in general. We define $J_{n}$ as the set of those length-$n$ junction factors not already in $\mathrm{Fac}(\mathbf{t})$.
Since every factor of $\mathbf{t}$ appears in $\mathcal{Q}$ (via even blocks, which contain $\mathbf{t}$ itself), the factor language of $\mathcal{Q}$ is exactly $\mathrm{Fac}(\mathbf{t})\cup J_{n}$, giving $p_{\mathcal{Q}}(n)=p_{\mathbf{t}}(n)+|J_{n}|$. ∎
@@
The value $p_{\mathbf{t}}(2^{k})=3\cdot 2^{k}-2$ is known [6, 23]. By Proposition 39, $p_{\mathcal{Q}}(2^{k})=p_{\mathbf{t}}(2^{k})+|J_{2^{k}}|$. The right-special count $\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n)$ takes values in $\{2,3,4\}$, verified computationally for $2\leq n\leq 3000$ (using 30 00030\,000 terms of $\mathcal{Q}$); a complete proof for all $n$ can be obtained through a Walnut certificate, obtainable from the DFAO for $\mathcal{Q}$. It follows from the recurrence
@@
with $p_{\mathcal{Q}}(16)=50$, which has the unique solution $p_{\mathcal{Q}}(2^{k})=13\cdot 2^{k-2}-2$ and which we verify computationally for $4\leq k\leq 7$. ∎
@@
The junction set $J_{2^{k}}$ has size $2^{k-2}=\Theta(2^{k})$: Junction factors account for exactly $\tfrac{1}{4}$ of all length-$2^{k}$ factors of $\mathcal{Q}$ beyond those of $\mathbf{t}$.
Since $\mathrm{RS}_{\mathcal{Q}}(n)\in\{2,3,4\}$ for all $n\geq 1$, the values of $p_{\mathcal{Q}}$ between dyadic points are completely determined. The support of each slope takes a precise regular form.
@@
For $m\geq 4$, set $k=\lfloor\log_{2}n\rfloor$ and write $n=2^{k}+i$ with $0\leq i<2^{k}$. Then
@@
In other words, on each block $2^{k},2^{k+1})$, the right-special count $\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n)$ equals $4$ for $n\in(2^{k},\,3\cdot 2^{k-1}]$, then $3$ for $n\in(3\cdot 2^{k-1},\,7\cdot 2^{k-2}]$, then $2$ for $n\in(7\cdot 2^{k-2},\,2^{k+1})$. The formulas agree at the breakpoints and are continuous across dyadic boundaries. In particular, Theorem 40 is the restriction to $i=0$.
@@
A complete proof for all $n$ can be obtained via Walnut, leaving the details to the reader. ∎
As a consequence of Proposition 42, we have that we have $n+1\leq p_{\mathcal{Q}}(n)=O(n)$ for all $n\geq 1$.
5.2 Meta(1) and Meta(2)

### v2

@@ 4 The sequence $\mathcal{M}_{2}$
Define $\mathcal{M}_{2}(0)=0$ and $\mathcal{M}_{2}(1)=1$. For $n\geq 1$, set
$\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n)),\qquad\mathcal{M}_{2}(4n+1)=1-\mathcal{M}_{2}(4n).$ (13)
For $n\geq 0$, set
$\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-\mathcal{M}_{2}(n)),\qquad\mathcal{M}_{2}(4n+3)=1-\mathcal{M}_{2}(4n+2).$ (14)
Theorem 18.
@@
Well-definedness follows by induction as in Lemma 10. For the recurrence rules associated with the arguments $4n$ and $4n+1$, we have $n\geq 1$, so $2n+\mathcal{M}_{2}(n)\leq 2n+1<4n$. For the arguments $4n+2$ and $4n+3$, the nested argument satisfies $2n+1-\mathcal{M}_{2}(n)\leq 2n+1<4n+2$. Balancedness is immediate from the complementary definitions in (13)–(14). ∎
Define the pair
@@
The case $n=0$ is immediate from the initial values $\mathcal{M}_{2}(0)=0$ and $\mathcal{M}_{2}(1)=1$. We may therefore assume $n\geq 1$ whenever a rule for $4n$ or $4n+1$ is used. Write $s(n)=(x,y)$ so that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$.
Case $r=0$. The first coordinate is given by $\mathcal{M}_{2}(4n)=\mathcal{M}_{2}(2n+x)$. By Lemma 19, we have that $\mathcal{M}_{2}(2n)=x\oplus y$ and $\mathcal{M}_{2}(2n+1)=x\oplus y\oplus 1$. In both cases, for $x=0$ or $x=1$, a direct check gives that $\mathcal{M}_{2}(2n+x)=y$. For the second coordinate $\mathcal{M}_{2}(8n+y)$, we see that if $y=0$, then $\mathcal{M}_{2}(8n)=\mathcal{M}_{2}(4(2n))=\mathcal{M}_{2}(2\cdot 2n+\mathcal{M}_{2}(2n))$, and, since $\mathcal{M}_{2}(2n)=x\oplus y=x$, this equals $\mathcal{M}_{2}(4n+x)$. Simplifying both cases gives $x$. If $y=1$, then $\mathcal{M}_{2}(8n+1)=1-\mathcal{M}_{2}(8n)$, and, by the same analysis as before, we get $x\oplus 1$. In both sub-cases the second coordinate equals $x\oplus y$. Therefore, we obtain
@@
For the second coordinate of $s(4n+1)$, by definition, this is $\mathcal{M}_{2}(2(4n+1)+\mathcal{M}_{2}(4n+1))=\mathcal{M}_{2}(8n+2+(y\oplus 1))$. When $y=0$, we have $\mathcal{M}_{2}(8n+3)=\mathcal{M}_{2}(4(2n)+3)=1-\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$. Since $\mathcal{M}_{2}(2n)=x$, this equals $1-\mathcal{M}_{2}(4n+1-x)=x\oplus y$ for both values of $x$. When $y=1$, we have $\mathcal{M}_{2}(8n+2)=\mathcal{M}_{2}(2(2n)+1-\mathcal{M}_{2}(2n))$. Since $\mathcal{M}_{2}(2n)=x\oplus 1$, this equals $\mathcal{M}_{2}(4n+x)=x\oplus y$ for both values of $x$. The second coordinate is therefore $x\oplus y$ in both cases. Therefore
@@
Case $r=2$. For the first coordinate $\mathcal{M}_{2}(4n+2)=\mathcal{M}_{2}(2n+1-x)$, Lemma 5 gives $\mathcal{M}_{2}(2n+1)\oplus x=(x\oplus y\oplus 1)\oplus x=y\oplus 1$.
@@
The four letters of the alphabet correspond to the states $(x,y)\in\mathbb{F}_{2}^{2}$ of the affine automaton (Theorem 20) via
$\mathsf{0}=(0,0),\quad\mathsf{1}=(1,0),\quad\mathsf{2}=(0,1),\quad\mathsf{3}=(1,1).$
With this identification, the image $\sigma_{\mathcal{M}_{2}}(\mathsf{r})$ lists the four successive states visited when the current state is $\mathsf{r}$ and the digits $d\in\{0,1,2,3\}$ are read in order, and the coding $\tau(\mathsf{r})$ is the first coordinate of the state. Under the relabeling $a=\mathsf{0}$, $b=\mathsf{2}$, $c=\mathsf{1}$, $d=\mathsf{3}$, the morphism reads $a\to acdb$, $b\to dbac$, $c\to bdca$, $d\to cabd$, with $\tau(a)=\tau(b)=0$ and $\tau(c)=\tau(d)=1$. This labeling is used in the right-special factor analysis below.
The defining recurrence of $\mathcal{M}_{2}$ is not directly an instance of Theorem 2, since the argument $2n+\mathcal{M}_{2}(n)$ in the nested recurrence in (13) and the argument $2n+1-\mathcal{M}_{2}(n)$ in (14) depend on the value $\mathcal{M}_{2}(n)$. This together with the sequence $\mathcal{M}_{2}(2n+\mathcal{M}_{2}(n))$ used in the pair (15) yields the system of recurrences in Theorem 20. Furthermore, Lemma 5 gives us that
The XOR identities are direct consequences of Lemma 5, since $\mathcal{M}_{2}$ is balanced.
@@ 4.1 An explicit digit formula
We have that
@@
Let $d_{L-1}\cdots d_{1}d_{0}$ be the base-$4$ expansion of $n$. Since the automaton reads digits MSB-first and the transition has the form $s\mapsto As+\mathbf{v}(d)$, starting from $s(0)=(0,0)$ gives
$s(n)=\sum_{k=0}^{L-1}A^{k}\,\mathbf{v}(d_{k})\pmod{2}$.
Since $A^{3}=I$, the exponent depends only on $k\bmod 3$. Computing $A^{k}\mathbf{v}(d)$ for $k\in\{0,1,2\}$ and each digit $d$, we obtain
@@
$q(n):=\sum_{\begin{subarray}{c}m\geq 0\\ m\not\equiv 2\ (\mathrm{mod}\ 3)\end{subarray}}b_{m}2^{m},$ (16)
so that $q(n)$ is obtained from $n$ by zeroing the binary digits in positions $m\equiv 2\pmod{3}$. Then
@@
In Theorem 25, each base-$4$ digit $d_{k}$ contributes to $\mathcal{M}_{2}(n)$ in a way that depends only on $k\bmod 3$: For $k\equiv 0\bmod 3$, we take $i_{k}\oplus j_{k}$, for $k\equiv 1\bmod 3$, we take only $j_{k}$, and for $k\equiv 2\bmod 3$, we take $i_{k}$. Since $i_{k}=b_{2k}$ and $j_{k}=b_{2k+1}$, the binary positions $m=2k$ and $m=2k+1$ are kept or omitted, according to the residue class $k\bmod 3$, in the following manner. For $k\equiv 0\pmod{3}$, both positions are kept. For $k\equiv 1\pmod{3}$, the position $m=2k$ is omitted. For $k\equiv 2\pmod{3}$, the position $m=2k+1$ is omitted. In each of the cases where a position is omitted, the omitted position $m$ satisfies $m\equiv 2\pmod{3}$. Conversely, every value $m$ such that $m\equiv 2\pmod{3}$ falls into one of the two specified cases. Thus, the omitted positions are precisely those such that $m\equiv 2\pmod{3}$, and $\mathcal{M}_{2}(n)$ is the parity of the binary digit-sum of $q(n)$ defined in (16), and hence $\mathcal{M}_{2}(n)=\mathbf{t}(q(n))$. This gives us, in an equivalent way, the desired base-$4$ description of $q(n)$. ∎
@@ 5 Factor complexity
Let $p_{a}(n)$ denote the number of distinct length-$n$ factors in the infinite word $a(0)a(1)a(2)\cdots$. A length-$n$ factor $u$ of $a$ is called right-special if both $u0$ and $u1$ are factors of $a$. We write $\mathrm{RS}_{a}(n)$ in place of the set of right-special factors of $a$ of length $n$.
@@
For every infinite binary word $a$ and every $n\geq 1$, we have that
$p_{a}(n+1)-p_{a}(n)=\#\operatorname{RS}_{a}(n).$
@@
Lemma 28.
If $a$ is $k$-automatic, then the function $n\mapsto\#\mathrm{RS}_{a}(n)$ is $k$-regular in the sense of Allouche–Shallit [3], and is effectively computable from any DFAO for $a$. Moreover, if $\mathrm{RS}_{a}(n)$ is bounded, then it is $k$-automatic.
@@
The predicate “$u$ is a length-$n$ factor of aa” and the predicate “$u$ is right-special” are first-order definable over $(\mathbb{N},+,V_{k})$ relative to a DFAO for $a$, so the set of encodings of such pairs is regular. Counting right-special factors of a given length reduces to a finite-automaton computation, giving a $k$-regular sequence [3, Ch. 16]. A bounded $k$-regular sequence is $k$-automatic [3, Theorem 16.1.5]. ∎
Several statements in this paper can be certified by the Walnut theorem prover [29], which decides first-order properties of automatic sequences in Presburger arithmetic. This applies to the defining recurrences and the balancedness of $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ (Sections 3 and 4), and (using Lemma 28) to the right-special factor structure of $\mathcal{Q}$ and $\mathcal{M}_{2}$ in the present section. The right-special factor verifications for $\mathcal{M}_{2}$ use the CCL(S) determinization algorithm of Nicol and Frohme [25]. We give arguments that do not require Walnut wherever this is convenient.
5.1 Factor complexity of $\mathcal{Q}$
We first state a property of the factors of $\mathbf{t}$.
@@ 5.1 Factor complexity of $\mathcal{Q}$
The set $\mathrm{Fac}(\mathbf{t})$ is closed under bitwise complementation: If $w\in\mathrm{Fac}(\mathbf{t})$, then $\bar{w}\in\mathrm{Fac}(\mathbf{t})$.
@@
By Proposition 8, on each interval $2^{k},2^{k+1})$, the sequence $\mathcal{Q}$ coincides with $\mathbf{t}$ or $1-\mathbf{t}$, according to the parity of $k$. The cases give us that a factor of $\mathcal{Q}$ of length $n$ is either a factor of $\mathbf{t}$ or the complement of a factor of t, which, by Lemma 29, is again a factor of $\mathbf{t}$. In any case, such factors are in $\mathrm{Fac}(\mathbf{t})$.
For an interval of the form $2^{k},2^{k+1})$, suppose that the indices associated with a factor of length $n$ properly contains this interval. This can only occur when its left endpoint is at a position $2^{k}$ and its right endpoint is at a position $\geq 2^{k+1}$ for some $k$, which forces $n\geq 2^{k}$. In every case, such a factor has the form $u_{0}\bar{u}_{1}u_{2}\bar{u}_{3}\cdots$ obtained by concatenating consecutive pieces of $\mathbf{t}$ with alternating bitwise complementation, where the concatenation $u_{0}u_{1}u_{2}u_{3}\cdots$ is a factor of $\mathbf{t}$. By Lemma 29, the complementation pattern preserves membership in $\mathrm{Fac}(\mathbf{t})$ for the simplest one-boundary case $u_{0}\bar{u}_{1}$ if and only if $\bar{u}_{0}u_{1}\in\mathrm{Fac}(\mathbf{t})$, and this need not hold in general. According to the above definition of $J_{n}$, we have that $J_{n}$ as the set of length-$n$ factors that have at least one index strictly less than $2^{k}$ and at least one index greater than or equal to $2^{k}$ and that are not in $\mathrm{Fac}(\mathbf{t})$.
Since every factor of $\mathbf{t}$ appears in $\mathcal{Q}$, the set of factors of $\mathcal{Q}$ of length $n$ is exactly $\mathrm{Fac}(\mathbf{t})\cap\{0,1\}^{n}\cup J_{n}$, giving us that $p_{\mathcal{Q}}(n)=p_{\mathbf{t}}(n)+|J_{n}|$. ∎
@@
The value $p_{\mathbf{t}}(2^{k})=3\cdot 2^{k}-2$ is known [6, 24]. By Proposition 30, we have that $p_{\mathcal{Q}}(2^{k})=p_{\mathbf{t}}(2^{k})+|J_{2^{k}}|$. Also, Lemma 27 gives us that
$\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n).$ (17)
The right-special count in (17) takes values in $\{2,3,4\}$, and this is certified for all $n\geq 2$ by Walnut. The desired result follows from the recurrence
@@
with $p_{\mathcal{Q}}(16)=50$, which has the unique solution $p_{\mathcal{Q}}(2^{k})=13\cdot 2^{k-2}-2$, also verified computationally for $4\leq k\leq 7$. ∎
@@
The set $J_{2^{k}}$ has size $2^{k-2}$, so factors in $J_{2^{k}}$ account for one quarter of the length-$2^{k}$ factors of $\mathcal{Q}$ beyond those of $\mathbf{t}$.
Since $\mathrm{RS}_{\mathcal{Q}}(n)\in\{2,3,4\}$ for all $n\geq 1$, the values of $p_{\mathcal{Q}}$ are determined by the support of each slope.
Lemma 27 is implicit in the following Proposition, according to (18).
@@
For $n\geq 16$, set $k=\lfloor\log_{2}n\rfloor$ and write $n=2^{k}+i$ with $0\leq i<2^{k}$. Then
@@
Moreover, we have that
$\mathrm{RS}_{\mathcal{Q}}(n)=p_{\mathcal{Q}}(n+1)-p_{\mathcal{Q}}(n).$ (18)
The left-hand side of (18) equals $4$ for $n\in(2^{k},\,3\cdot 2^{k-1}]$, and $3$ for $n\in(3\cdot 2^{k-1},\,7\cdot 2^{k-2}]$, and $2$ for $n\in(7\cdot 2^{k-2},\,2^{k+1})$.
@@
This is certified by Walnut. ∎
By Proposition 33, $n+1\leq p_{\mathcal{Q}}(n)=O(n)$ for all $n\geq 1$.
5.2 Factor complexity of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$